OpenStudy (anonymous):

factor completely 3s^2-30s+75

OpenStudy (anonymous):

GCF of 3\[=3(s^2-10s+25)\] perfect square trinomial since -5*(-5)=25 and -5-5=-10\[=3(s-5)^2\]this is the factored form